OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) - 4. Token and fan?!?!

OpenStudy (aum):

3(x + 1) = -2(x - 1) - 4 Distribute the 3 and the -2 3x + 3 = -2x + 2 - 4 3x + 3 = -2x - 2 add 2x to both sides: 5x + 3 = -2 subtract 3 from both sides: 5x = -5 divide both sides by 5: x = -1

OpenStudy (anonymous):

Thank you so much! Can you help me with 1 more question?

OpenStudy (aum):

I can try.

OpenStudy (anonymous):

Solve for x: 4 over 5 x + 4 over 3 = 2x x = __________________ Write your answer as a fraction in simplest form. Use the "/" symbol for the fraction bar.

OpenStudy (aum):

Problem not clear. Is it 4 / (5x + 4) / 3 = 2x ?

OpenStudy (aum):

\( \Large \frac{4}{5x + 4/3} = 2x\) ?

OpenStudy (anonymous):

Doesn't show any parenthesis

OpenStudy (anonymous):

I'll type how it is on the question

OpenStudy (anonymous):

\[\frac{4 }{ 5 }+\frac{ 4 }{ 3 }=2\]

OpenStudy (anonymous):

Sorry, first time using the equations on this site

OpenStudy (anonymous):

Oops, forgot to put the x to the right of 4/5

OpenStudy (anonymous):

and by the x :P

OpenStudy (anonymous):

2*

OpenStudy (aum):

\(\Large \frac 45x + \frac 43 = 2x\) Multiply throughout by 15: \(\Large 15*\frac 45x + 15*\frac 43 = 15*2x\) 12x + 20 = 30x subtract 12x 20 = 18x divide by 18 20/18 = x x = 20/18 = 10/9

OpenStudy (anonymous):

Thank you.

OpenStudy (aum):

You are welcome.